# Hilbert property for low-genus families and degree-one del Pezzo surfaces

### v1

@@ 3. Low-genus families on degree-one del Pezzo surfaces
The first property follows from injectivity of specialization on the Picard group (see e.g. [29, §5]). More precisely, in (2) we mean that there exists a finite collection of irreducible covers $\pi_{i}:Y_{i}\to U,i=1,\ldots,r$ of finite degrees >1>1 such that $\operatorname{rk}\operatorname{Pic}X_{u}>\operatorname{rk}\operatorname{Pic}X_{\eta}$ if and only if $u\in\operatorname{im}(Y_{i}(\kappa(u))\to U(\kappa(u)))$. This holds by applying Hilbert’s Irreducibility Theorem to the family of degree-240 polynomials whose roots parametrize the lines in each surface in our family. The Galois orbits of lines correspond to the irreducible factors of these polynomials and determine the Picard rank. The Picard rank can only increase if an irreducible factor becomes reducible upon specialization, which occurs over a thin locus. By a result of Kloosterman [16, Lem. 2.5, Prop. 2.9], reproving a result of Desjardins and Naskrecki [10, Sec. 4.2]), the del Pezzo surface

### v2

@@ 3. Low-genus families on degree-one del Pezzo surfaces
The first property follows from injectivity of specialization on the Picard group (see e.g. [34, §5]). More precisely, in (2) we mean that there exists a finite collection of irreducible covers $\pi_{i}:Y_{i}\to U,i=1,\ldots,r$ of finite degrees >1>1 such that $\operatorname{rk}\operatorname{Pic}X_{u}>\operatorname{rk}\operatorname{Pic}X_{\eta}$ if and only if $u\in\operatorname{im}(Y_{i}(\kappa(u))\to U(\kappa(u)))$. This holds by applying Hilbert’s Irreducibility Theorem to the family of degree-240 polynomials whose roots parametrize the lines in each surface in our family. The Galois orbits of lines correspond to the irreducible factors of these polynomials and determine the Picard rank. The Picard rank can only increase if an irreducible factor becomes reducible upon specialization, which occurs over a thin locus. Thus the Picard rank can only increase over a thin locus. By a result of Kloosterman [20, Lem. 2.5, Prop. 2.9], reproving a result of Desjardins and Naskrecki [12, Sec. 4.2]), the del Pezzo surface